i have no idea. It says evaluate the exonential expression. write answers as a fraction in simplest form

1. (1/2)1-


2. 73 with 0 power


3. 13-x=



4. 4-2


i don't know what this stuff is

Answer by tutorcecilia(2144)

1. (1/2)-1 = 1/2^-1 = 1/(1/2)^1 = 2
When a a number is raised to a negative exponent, turn it into a positive exponent by taking the inverse of the number. To get the inverse, just flip the number and place it under or ontop of 1 (depending on the problem).
.
Example: 3^(-2) = Flip it = 1/3^2
The inverse of 3 is 1/3. Be sure to include the positive exponent.
.
Example 1/3^-2 = Flip it = 3^2.
.
2. 73 with 0 power = 73^0 = 0
Any number raised to zero equals zero.
Example: 3^0 = 0
.
.
3. 13-x= 13^(-x) = Flip it = 1/(13^x)
.
.
4. 4-2= 4^(-2) = Flip it = 1/4^(2) = 1/16